Problem 1: Fenoprofen persones d PK, 4.5 What is percent ionized at pH 7.5? CH-COOH D CH3 ES 131
Answer #1

Ch - coold CH3 pka = 4.5 - logka = 4.5 Ra - 10-4.5 = 316416-5 Percrent inital PH = 75 - loy (H) = 75 *(1+0) = 20 -75 [H] = 3,